I know the answer to X is 0, but I can't figure out how to get that answer. The problem is:
1/2 (x-3)+ (3/2-x)=5x

Answer by Alan3354(69443)

1/2 (x-3)+ (3/2-x)=5x
------------
Multiply by 2
(x-3) + (3-2x) = 10x
-x = 10x
Only 0 will satisfy that.
Otherwise -1 = 10
It's an example of an equation someone dreamed up that has no application in reality.